# TALEC: Teach Your LLM to Evaluate in Specific Domain with In-house Criteria by Criteria Division and Zero-shot Plus Few-shot

### v1

@@ Abstract
*Kaiqi Zhang1,2 *Shuai Yuan1 *Honghan Zhao1 1ByteDance Inc 2University of Electronic Science and Technology of China {zhangkaiqi.zlkqz, yuanshuai.irbaozi, zhaohonghan}@bytedance.com
@@ 1 Introduction
In the end of this paper, we compare TALEC with the other automatic evaluation method and humans. TALEC demonstrates a strong capability to accurately reflect human preferences and achieves a correlation of over 80% with human judgments, outperforming many other methods and even the inter-human correlation in some tasks
@@ 3.1 Customized Business Evaluation Criteria
Label Name | Description | Score |
Not Meeting the Requirements | Answer failed to strictly enforce the requirements. | 0 |
Incorrect Answer/Unrelated Matching Results | Containing incorrect information or mismatched responses. | 0 |
Refusal to Answer | The model explicitly shows a refusal to answer when it can. | 0 |
Untranslated Text | Large portions of non-Chinese language in responses. | 0 |
Confusing Answers | Answers contain messy code or content that interferes with reading. | 0 |
External Links or Diversions | External links or obvious diversionary behavior in the answer. | 0 |
Stiffness | The response lacked anthropomorphic expression. | 1 |
Repetitive Expression | Repeated expressions in the answer. | 1 |
Subject Imprecision | Redundant content or unclear subject in the response | 1 |
Incomplete Answers | Failure to address all needs in the directive | 1 |
@@ 4.1 Criteria Division
Method | Task | Spearman | |
eval | test | ||
Standard Prompt Paradigm (Division) | Sentiment Analysis | 0.9579 | 0.9693 |
Knowledge QA | 0.4945 | 0.5063 | |
Search QA | 0.8263 | 0.8487 | |
Title Generation | 0.9207 | 0.9006 | |
Non-division | Sentiment Analysis | 0.3735 | 0.2905 |
Knowledge QA | 0.5398 | 0.4251 | |
Search QA | 0.689 | 0.4691 | |
Title Generation | 0.3763 | 0.4296 |
Method | task | label | Acc | F1/Precision/Recall | ||
eval | test | eval | test | |||
Standard Prompt Paradigm (Division) | Knowledge QA | Incorrect Answer /Unrelated Matching Results | 0.8298 | 0.7737 | 0.52/0.3611/0.9286 | 0.2439/0.1667/0.4545 |
Non-division | Knowledge QA | Incorrect Answer /Unrelated Matching Results | 0.9149 | 0.8321 | 0.5714/0.5714/0.5714 | 0.1481/0.125/0.1818 |
@@ 4.2 Combine Zero-shot with Few-shot
Method | Task | Spearman | |
eval | test | ||
Standard Prompt Paradigm (Single-turn wo Zero-shot) | Sentiment Analysis | 0.9579 | 0.9693 |
Knowledge QA | 0.4945 | 0.5063 | |
Search QA | 0.8263 | 0.8487 | |
Title Generation | 0.9207 | 0.9006 | |
Multi-turn with Zero-shot | Sentiment Analysis | 0.9536 | 0.9485 |
Knowledge QA | 0.8597 | 0.7089 | |
Search QA | 0.8574 | 0.9438 | |
Title Generation | 0.915 | 0.9206 |
Method | task | label | Acc | F1/Precision/Recall | ||
eval | test | eval | test | |||
Standard Prompt Paradigm (Single-turn wo Zero-shot) | Knowledge QA | Incorrect Answer /Unrelated Matching Results | 0.8298 | 0.7737 | 0.52/0.3611/0.9286 | 0.2439/0.1667/0.4545 |
Multi-turn with Zero-shot | Knowledge QA | Incorrect Answer /Unrelated Matching Results | 0.9645 | 0.9051 | 0.8148/0.8462/0.7857 | 0.5185/0.4375/0.6364 |
@@ 4.3 SFT vs ICL
Method | Task | Spearman | |
eval | test | ||
Standard Prompt Paradigm (GPT4 + ICL) | Sentiment Analysis | 0.9579 | 0.9693 |
Knowledge QA | 0.4945 | 0.5063 | |
Search QA | 0.8263 | 0.8487 | |
Title Generation | 0.9207 | 0.9006 | |
Qwen-72B-Chat + ICL | Sentiment Analysis | 0.7676 | 0.7578 |
Knowledge QA | 0.2969 | 0.1766 | |
Search QA | 0.519 | 0.4513 | |
Title Generation | 0.1585 | 0.1392 | |
Qwen-72B-Chat + SFT | Sentiment Analysis | 0.4565 | 0.4211 |
Knowledge QA | 0.1886 | 0.2444 | |
Search QA | 0.5705 | 0.4772 | |
Title Generation | 0.2738 | 0.2671 |
@@ 5.1 Repeat Descriptions of Evaluation Criteria
Method | Task | Spearman | |
eval | test | ||
Standard Prompt Paradigm (Repeat descriptions) | Sentiment Analysis | 0.9579 | 0.9693 |
Knowledge QA | 0.4945 | 0.5063 | |
Search QA | 0.8263 | 0.8487 | |
Title Generation | 0.9207 | 0.9006 | |
Non-repetition | Sentiment Analysis | 0.9553 | 0.9658 |
Knowledge QA | 0.4911 | 0.4823 | |
Search QA | 0.8208 | 0.838 | |
Title Generation | 0.6618 | 0.6815 |
@@ 5.2 Standardize the Format of Examples
Method | Task | Spearman | |
eval | test | ||
Arbitrariness | Sentiment Analysis | 0.9551 | 0.9515 |
Knowledge QA | 0.5242 | 0.5459 | |
Search QA | 0.7946 | 0.814 | |
Title Generation | 0.535 | 0.5729 | |
Standard (Non-repetition) | Sentiment Analysis | 0.9553 | 0.9658 |
Knowledge QA | 0.4911 | 0.4823 | |
Search QA | 0.8208 | 0.838 | |
Title Generation | 0.6618 | 0.6815 |
@@ 6.2 Comparison with Human Annotation
Task | Spearman |
Sentiment Analysis | 0.9054 |
Knowledge QA | 0.6523 |
Search QA | 0.7973 |
Title Generation | 0.8772 |

### v2

@@ Abstract
*Kaiqi Zhang1,2 *Shuai Yuan1 *Honghan Zhao11ByteDance China 2University of Electronic Science and Technology of Chinazhangkaiqi.zlk@gmail.com{yuanshuai.irbaozi, zhaohonghan}@bytedance.com
00footnotetext: First submitted on June 25, 2024.
@@ 1 Introduction
In the end of this paper, we compare TALEC with the other automatic evaluation method and humans. TALEC demonstrates a strong capability to accurately reflect human preferences and
achieves a correlation of over 80% with human judgments, outperforming many other methods and even the inter-human correlation in some tasks
Figure 1: The overall process of TALEC
@@ 3.1 Customized Business Evaluation Criteria
Label Name
Description
Score
Not Meeting the Requirements
Answer failed to strictly enforce the requirements.
0
Incorrect Answer/Unrelated Matching Results
Containing incorrect information or mismatched responses.
0
Refusal to Answer
The model explicitly shows a refusal to answer when it can.
0
Untranslated Text
Large portions of non-Chinese language in responses.
0
Confusing Answers
Answers contain messy code or content that interferes with reading.
0
External Links or Diversions
External links or obvious diversionary behavior in the answer.
0
Stiffness
The response lacked anthropomorphic expression.
1
Repetitive Expression
Repeated expressions in the answer.
1
Subject Imprecision
Redundant content or unclear subject in the response
1
Incomplete Answers
Failure to address all needs in the directive
1
Table 1: Labels to be evaluated. The labels can be divided into two categories: acceptable labels(score=1) and unacceptable labels(score=0). If any unacceptable label appears, the final sore will be 0. If unacceptable label does not appear but any acceptable label appears, the final sore will be 1. If no labels appears, the final sore will be 2 (full score). Note that we don’t use our model-based method to judge the word count because LLM can’t accurately count the number of words.
@@ 4.1 Criteria Division
Method
Task
Spearman
eval
test
Standard PromptParadigm(Division)
Sentiment Analysis
0.9579
0.9693
Knowledge QA
0.4945
0.5063
Search QA
0.8263
0.8487
Title Generation
0.9207
0.9006
Non-division
Sentiment Analysis
0.3735
0.2905
Knowledge QA
0.5398
0.4251
Search QA
0.689
0.4691
Title Generation
0.3763
0.4296
Table 2: Comparison between criteria division and non-division. The details can be found in Table 11 and Table 15.
Method
task
label
Acc
F1/Precision/Recall
eval
test
eval
test
Standard PromptParadigm(Division)
KnowledgeQA
Incorrect Answer/UnrelatedMatching Results
0.8298
0.7737
0.52/0.3611/0.9286
0.2439/0.1667/0.4545
Non-division
KnowledgeQA
Incorrect Answer/UnrelatedMatching Results
0.9149
0.8321
0.5714/0.5714/0.5714
0.1481/0.125/0.1818
Table 3: The detailed score comparison on "Incorrect Answer/Unrelated Matching Results" label in Knowledge QA task. The experimental setups are the same as in Table 2.
@@ 4.2 Combine Zero-shot with Few-shot
Method
Task
Spearman
eval
test
Standard PromptParadigm(Single-turnwo Zero-shot)
Sentiment Analysis
0.9579
0.9693
Knowledge QA
0.4945
0.5063
Search QA
0.8263
0.8487
Title Generation
0.9207
0.9006
Multi-turnwith Zero-shot
Sentiment Analysis
0.9536
0.9485
Knowledge QA
0.8597
0.7089
Search QA
0.8574
0.9438
Title Generation
0.915
0.9206
Table 4: Comparison between the results of single-turn without zero-shot and multi-turn with zero-shot. The details can be found in Table 11 and Table 12.
Method
task
label
Acc
F1/Precision/Recall
eval
test
eval
test
Standard PromptParadigm(Single-turn wo Zero-shot)
KnowledgeQA
Incorrect Answer/UnrelatedMatching Results
0.8298
0.7737
0.52/0.3611/0.9286
0.2439/0.1667/0.4545
Multi-turn with Zero-shot
KnowledgeQA
Incorrect Answer/UnrelatedMatching Results
0.9645
0.9051
0.8148/0.8462/0.7857
0.5185/0.4375/0.6364
Table 5: The detailed score comparison on Incorrect Answer/Unrelated Matching Results label in Knowledge QA task. The experimental setups are the same as in Table 4.
As you can see in Table 4, multi-turn with zero-shot approach can improve scores on all task. Especially in "Incorrect Answer/Unrelated Matching Results" label of Knowledge QA, Table 5 shows greatly improvement in this label.
@@ 4.3 SFT vs ICL
Method
Task
Spearman
eval
test
Standard Prompt Paradigm(GPT4 + ICL)
Sentiment Analysis
0.9579
0.9693
Knowledge QA
0.4945
0.5063
Search QA
0.8263
0.8487
Title Generation
0.9207
0.9006
Qwen-72B-Chat + ICL
Sentiment Analysis
0.7676
0.7578
Knowledge QA
0.2969
0.1766
Search QA
0.519
0.4513
Title Generation
0.1585
0.1392
Qwen-72B-Chat + SFT
Sentiment Analysis
0.4565
0.4211
Knowledge QA
0.1886
0.2444
Search QA
0.5705
0.4772
Title Generation
0.2738
0.2671
Table 6: Comparison between GPT4 with in-context learning, Qwen-72B-Chat with in-context learning and Qwen-72B-Chat SFT without in-context learning.The details can be found in Table 11, Table 13 and Table 14.
@@ 5.1 Repeat Descriptions of Evaluation Criteria
Figure 2: This figure illustrates an instance where the model fails to consider the descriptions within the System Prompt, juxtaposed with a contrasting example where the model redundantly repeats the descriptions before proceeding to evaluation.
@@
Method
Task
Spearman
eval
test
Standard PromptParadigm(Repeat descriptions)
Sentiment Analysis
0.9579
0.9693
Knowledge QA
0.4945
0.5063
Search QA
0.8263
0.8487
Title Generation
0.9207
0.9006
Non-repetition
Sentiment Analysis
0.9553
0.9658
Knowledge QA
0.4911
0.4823
Search QA
0.8208
0.838
Title Generation
0.6618
0.6815
Table 7: Comparison of the experimental results for repeating descriptions and non-repetition. The details can be found in Table 11 and Table 16.
@@ 5.2 Standardize the Format of Examples
Method
Task
Spearman
eval
test
Arbitrariness
Sentiment Analysis
0.9551
0.9515
Knowledge QA
0.5242
0.5459
Search QA
0.7946
0.814
Title Generation
0.535
0.5729
Standard(Non-repetition)
Sentiment Analysis
0.9553
0.9658
Knowledge QA
0.4911
0.4823
Search QA
0.8208
0.838
Title Generation
0.6618
0.6815
Table 8: Comparison between the results of few-shot with arbitrary format and standard format. The details can be found in Table 10 and Table 16.
@@ 6.2 Comparison with Human Annotation
Task
Spearman
Sentiment Analysis
0.9054
Knowledge QA
0.6523
Search QA
0.7973
Title Generation
0.8772
Table 9: Alignment degree of the manual annotations before quality inspection.
@@ Appendix A Prompt Template
The prompt templates are exemplified in the Figure 3 and Figure 4.
Figure 3: The prompt format for evaluation tasks is depicted in the following sequence: the system-generated prompt, containing explicit task requirements and evaluation criteria explanations, aids the model in comprehending the assessment’s scope and complexity. This is succeeded by the few-shot section, which presents a variety of examples in the same format, including both positive and negative instances, enabling the model to adopt the sample structure for its output. Finally, the output produced by LLM, is the subject of evaluation.Figure 4: This figure illustrates an instance of the prompt format for Combining Zero-shot with Few-shot. We combine system Prompt, Shots, answer output by zero-shot, answer output by few-shot to get a new context and use this con-text to judge again. Then we get the final result.